OpenStudy (anonymous):

Move everything away from y by using the additive inverse/opposite 3x - y = 5

OpenStudy (tommo_lcfc):

3x - y = 5 5 - y = 3x y = 5 - 3x

OpenStudy (muhammad_nauman_umair):

3x-3x-y=5-3x -y=5-3x -1*-y=-1*(5-3x) y=3x-5